# Large Language Models can be Guided to Evade AI-Generated Text Detection

### v5

@@ Appendix I Examples of Text Generated by SICO
SICO-Para: The code’s layout was systematically arranged for user convenience, with the primary objective of arranging the control and graphic codes within proximal access to facilitate swift identification. The layout allotted the first two columns, constituting a total of 32 positions, solely for control codes, while the remaining 94 columns (94 positions) were allocated for graphic codes, promoting seamless recognition and expediting processing operations. Furthermore, the code’s malleability was considered a paramount attribute - the code was designed to be effortlessly modified, accommodating the addition and deletion of codes as per necessity. These inherent characteristics rendered the code an extremely pragmatic and adaptable tool for a wide range of programming and data processing applications.
0.00
@@
SICO-Gen: In order to facilitate easy and efficient use, the code was structured in a manner wherein control codes and graphic codes were distinctly grouped together, allowing for streamlined identification. This design allowed the first two columns (equivalent to 32 positions) to be exclusively reserved for control codes, ensuring that they were easily and rapidly located. The following columns, including columns three to seven, were allocated for graphic codes - providing ample space to incorporate them in a comprehensive manner. By distributing codes systematically and allocating ample space for each grouping, the code structure was organized for optimal use and readability. The efficiency of this coding scheme underscores its practical and logistical utility, while its unique design contributes to its widely accepted success and usage in various computer systems. The contextual background supplied enriches the readers’ understanding of the code structure’s significance and its potential impact on a system’s proficiency.
0.02
@@
Writing Start: Germany has a comprehensive school known as the Gesamtschule. While some German schools such as the Gymnasium and the Realschule have rather strict entrance requirements, the Gesamtschule does not have
$\mathcal{P}_{\text{AI}}$
@@
AI: Germany has a comprehensive school known as the Gesamtschule. While some German schools such as the Gymnasium and the Realschule have rather strict entrance requirements, the Gesamtschule does not have any such requirements. The Gesamtschule is a type of school that caters to students with varying academic abilities and backgrounds, and provides them with an equal opportunity for education. This school system aims to promote social integration and reduce educational inequality by offering a broad curriculum and flexible educational pathways. In addition, the Gesamtschule promotes the development of students’ individual strengths and talents through its inclusive and cooperative learning environment. Overall, the Gesamtschule represents a unique approach to education that prioritizes equal access and opportunity for all students.
0.99
@@
SICO-Para: Germany boasts a comprehensive educational institution known as the Gesamtschule. Unlike typical German schools such as the Gymnasium and Realschule, which impose strict entry prerequisites, the Gesamtschule embraces an inclusive admission criterion, ensuring equal education opportunities for students of diverse academic abilities and background. The Gesamtschule represents a collaboration between pupils with different skill sets paired with an extensive and elastic curriculum, thereby striving to promote social homogeneity and curb academic inequalities. Furthermore, the Gesamtschule’s inclusive and cooperative learning environment seeks to stimulate the cultivation of student strengths while also encouraging teamwork. Altogether, the Gesamtschule epitomizes an innovative educational frontier that prioritizes equality and accessibility for all students.
0.01
@@
SICO-Gen: An outstanding feature of the German education system is the Gesamtschule, which is a holistic or comprehensive school. In comparison to Gymnasium or Realschule where students are subject to strict requirements to gain admission, entrance to the Gesamtschule is less rigorous. This practice makes the school system more inclusive and allows students from diverse backgrounds to have access to quality education, fostering social unity. At Gesamtschule, students have the liberty to move between different academic tracks based on their demonstrated skills and interests, hence having a more personalized education experience. A blended curriculum is used; it includes theoretical and practical activities intended to prepare students for their future careers or academic pursuits. The Gesamtschule is a pioneering framework in the German schooling system because it promotes equal opportunities in education, educational diversity, and at the same time equips the students with skill sets vital for their future endeavors. The contextual background information provides a clear insight into the working of the Gesamtschule and underscores the importance of creating such models in fostering a more equitable and inclusive society.
0.00

### v6

@@ Appendix H Examples of Text Generated by SICO
SICO-Para: The code’s layout was systematically arranged for user convenience, with the primary objective of arranging the control and graphic codes within proximal access to facilitate swift identification. The layout allotted the first two columns, constituting a total of 32 positions, solely for control codes, while the remaining 94 columns (94 positions) were allocated for graphic codes, promoting seamless recognition and expediting processing operations. Furthermore, the code’s malleability was considered a paramount attribute - the code was designed to be effortlessly modified, accommodating the addition and deletion of codes as per necessity. These inherent characteristics rendered the code an extremely pragmatic and adaptable tool for a wide range of programming and data processing applications. 0.00
@@
SICO-Gen: In order to facilitate easy and efficient use, the code was structured in a manner wherein control codes and graphic codes were distinctly grouped together, allowing for streamlined identification. This design allowed the first two columns (equivalent to 32 positions) to be exclusively reserved for control codes, ensuring that they were easily and rapidly located. The following columns, including columns three to seven, were allocated for graphic codes - providing ample space to incorporate them in a comprehensive manner. By distributing codes systematically and allocating ample space for each grouping, the code structure was organized for optimal use and readability. The efficiency of this coding scheme underscores its practical and logistical utility, while its unique design contributes to its widely accepted success and usage in various computer systems. The contextual background supplied enriches the readers’ understanding of the code structure’s significance and its potential impact on a system’s proficiency. 0.02
@@
Writing Start: Germany has a comprehensive school known as the Gesamtschule. While some German schools such as the Gymnasium and the Realschule have rather strict entrance requirements, the Gesamtschule does not have $\mathcal{P}_{\text{AI}}$
@@
AI: Germany has a comprehensive school known as the Gesamtschule. While some German schools such as the Gymnasium and the Realschule have rather strict entrance requirements, the Gesamtschule does not have any such requirements. The Gesamtschule is a type of school that caters to students with varying academic abilities and backgrounds, and provides them with an equal opportunity for education. This school system aims to promote social integration and reduce educational inequality by offering a broad curriculum and flexible educational pathways. In addition, the Gesamtschule promotes the development of students’ individual strengths and talents through its inclusive and cooperative learning environment. Overall, the Gesamtschule represents a unique approach to education that prioritizes equal access and opportunity for all students. 0.99
@@
SICO-Para: Germany boasts a comprehensive educational institution known as the Gesamtschule. Unlike typical German schools such as the Gymnasium and Realschule, which impose strict entry prerequisites, the Gesamtschule embraces an inclusive admission criterion, ensuring equal education opportunities for students of diverse academic abilities and background. The Gesamtschule represents a collaboration between pupils with different skill sets paired with an extensive and elastic curriculum, thereby striving to promote social homogeneity and curb academic inequalities. Furthermore, the Gesamtschule’s inclusive and cooperative learning environment seeks to stimulate the cultivation of student strengths while also encouraging teamwork. Altogether, the Gesamtschule epitomizes an innovative educational frontier that prioritizes equality and accessibility for all students. 0.01
@@
SICO-Gen: An outstanding feature of the German education system is the Gesamtschule, which is a holistic or comprehensive school. In comparison to Gymnasium or Realschule where students are subject to strict requirements to gain admission, entrance to the Gesamtschule is less rigorous. This practice makes the school system more inclusive and allows students from diverse backgrounds to have access to quality education, fostering social unity. At Gesamtschule, students have the liberty to move between different academic tracks based on their demonstrated skills and interests, hence having a more personalized education experience. A blended curriculum is used; it includes theoretical and practical activities intended to prepare students for their future careers or academic pursuits. The Gesamtschule is a pioneering framework in the German schooling system because it promotes equal opportunities in education, educational diversity, and at the same time equips the students with skill sets vital for their future endeavors. The contextual background information provides a clear insight into the working of the Gesamtschule and underscores the importance of creating such models in fostering a more equitable and inclusive society. 0.00